Bipti is a chemist at Venkatanarasimharajuvaripeta Chemical Ltd. She has published her results of the isomers of 2-Fluoropropane . The resultant masses of the isomers chosen are as follow:

5.6 gr, 5.1 gr, 4.6gr, 4.8 gr, 5.7 gr and 6.4 gr. Construct a 90 % confidence level for these sample data.
Answer #1

Table for sample mean and sample standard deviation -

Xi (X_i - \bar{X})^{2}
5.6
5.1
4.6
4.8
5.7
6.4
0.0529
0.0729
0.5929
0.3249
0.1089
1.0609
\sum X_i = 32.2 \sum (X_i - \bar{X})^{2} = 2.213

Sample size = n = 6

Sample Mean = \bar{X} =\frac{\sum X_i}{n} = \frac{32.2}{6} = 5.37

Sample standard deviation =

S = \sqrt{\frac{\sum(X_i - \bar{X})^{2} }{n-1}} = \sqrt{\frac{2.213}{5}} = 0.665​​​​​​90% confidence interval is given by -

[\bar{x}\pm z_{0.10/2}*\frac{s}{\sqrt{n}}]

where, z_{0.10/2} is the Critical value of z for two tailed test at 0.10 level of significance = 1.65

(It can be obtained using the z table by finding the z corresponding to the area close to 0.05)

Hence, the confidence interval will be given by -

[5.37\pm 1.65*\frac{0.665}{\sqrt{6}}]

=[5.37\pm 0.45]

= [4.92, 5.81]